# Practitioner Motives to Use Different Hyperparameter Optimization Methods

### v3

@@ Abstract.
Programmatic hyperparameter optimization (HPO) methods, such as Bayesian optimization and evolutionary algorithms, show high sampling efficiency in finding optimal hyperparameter configurations in development of machine learning (ML) models. Yet, practitioners often use less sample-efficient HPO methods, such as grid search, which often results in under-optimized ML models. As a reason for this behavior, we suspect practitioners choose HPO methods based on different motives. Practitioner motives, however, still need to be clarified to enhance user-centered development of HPO tools. To uncover practitioner motives to use different HPO methods, we conducted 20 semi-structured interviews and an online questionnaire with 49 ML experts. By presenting main goals (e.g.,  increase ML model understanding) and contextual factors affecting practitioners’ selection of HPO methods (e.g., available computer resources), this study offers a conceptual foundation to better understand why practitioners use different HPO methods, supporting design of more user-centered and context-adaptive HPO tools in automated ML.
@@ 1. Introduction
Performance of machine learning (ML)models depends strongly on the choice of hyperparameter configurations (Bergstra and Bengio, 2012; Melis et al., 2018; Henderson et al., 2018; Chen et al., 2018; Zhang et al., 2021; Kadra et al., 2021). Finding optimal hyperparameter configurations for training ML models is, however, challenging—even for seasoned ML experts—because of large search spaces of hyperparameter values and commonly unknown relationships between ML model performance, hyperparameter configuration, and dataset. Consequently, practitioners—human developers of effective ML models—experiment with various hyperparameter configurations, often relying on trial and error to identify optimal configurations. The iterative process of exploring, testing, and adjusting hyperparameter configurations related to ML models is known as hyperparameter optimization (HPO).
@@
HPO by manual tuning is often cumbersome, tedious, and error-prone. To help practitioners with HPO, research with a technical focus (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)) developed several programmatic HPO methods, including grid search, random search, Bayesian optimization, and evolutionary algorithms, and implemented such methods as software tools (e.g., Hyperopt and Hyperband; (Bergstra et al., 2015; Li et al., 2018a)). The developed HPO methods, programmatic and non-programmatic ones, differ considerably in the way they optimize hyperparameter configurations. Such differences affect the suitability of HPO methods to satisfy requirements, such as reproducibility, high ML model performance, and high sample efficiency (Claesen and De Moor, 2015).
The superiority of HPO methods over others is shown based on conventional performance metrics from computer science, including minimization of generalization errors and increasing sample efficiency (Zöller and Huber, 2021; Gijsbers et al., 2022; Lindauer et al., 2022). Metrics related to ML model performance are commonly assumed to be sufficient to represent the principal motives of practitioners in HPO. This assumption is, however, questionable because practitioners often use HPO methods that are inferior according to conventional performance metrics. For example, practitioners often prefer to perform grid search over the more sample-efficient Bayesian optimization (Snoek et al., 2012).
@@
The dominant use of seemingly inferior HPO methods suggests that practitioners have additional motives for HPO. Such motives appear beyond fulfillment of conventional goals pursued in HPO, such as increasing ML model performance. Practitioner motives (i.e., goals influenced by contextual factors) to use HPO methods remain however unclear. The unclarity about why practitioners use which HPO methods inhibits more user-centered development of HPO methods and tools that support practitioners in attaining their goals, especially those beyond conventional performance metrics. To support development of more user-centered HPO methods and tools for automated ML (AutoML)(Vanschoren, 2019), practitioner motives for HPO need to be better understood. We approach the following research question: Why do practitioners choose different HPO methods?
@@
We applied a two-step research approach consisting of an interview study and a survey study based on an online questionnaire. First, we conducted semi-structured interviews with 20 ML experts to unveil HPO methods that practitioners commonly use, the goals pursued by practitioners when applying HPO methods, and the contextual factors that influence practitioners’ decisions for HPO methods to attain goals. Second, we performed an online survey with 49 participants to collect evidence for the external validity of the relevance of the HPO methods, goals, and contextual factors identified in the interviews.
@@
Our main ambition is to support more user-centric development of automated ML (AutoML) methods and tools for HPO by bridging the gap between technical advancements and human factors. By integrating practitioner motives into HPO research, we aim to ensure that future tools are not only technically sound but also aligned with real-world needs and decision-making processes. In particular, this work has three main contributions. First, this work presents a conceptual foundation comprised of principal goals (e.g., improving ML model performance and target audience compliance) and contextual factors (e.g., compute resources and method traceability) influencing selections of HPO methods. This is useful to enhance human-in-the-loop ML by clarifying information needs and improving practitioner engagement. Moreover, the conceptual foundation supports goal-driven development of HPO methods and tools and new benchmarks with a focus on human factors. Second, we present a mapping of HPO methods, goals, and contextual factors to explain practitioners’ decision-making. This mapping informs user-centered HPO design and adaptive automation features. Moreover, it highlights key input parameters for developing context-sensitive HPO tools. Third, we analyzed practitioners’ perceived success with different HPO methods, revealing strengths and areas for improvement. These findings support better decision-support systems for selecting HPO methods. Moreover, the reported success rates help identify limitations of HPO methods, guiding development of better-tailored and more effective HPO methods and tools.
@@ 2. Background and Related Work
Mainly driven by technical advancements, HPO is a vital research field with much potential to shape development of ML models (Hutter et al., 2018). To support a better understanding of the results of this study, we briefly describe the technical foundations of principal HPO methods. Moreover, we elucidate related research on involving human beings in AutoML, with a focus on automated HPO.
@@ 2.1. Hyperparameter Optimization Methods
There are five principal, model-agnostic HPO methods that are commonly used: manual tuning, grid search, random search, evolutionary algorithms, and Bayesian optimization. [^1] There are also gradient-based methods, but because these methods are often perceived as being brittle and are often model-specific, they are out of the scope of this study. Those HPO methods are briefly described in the following.
@@
Manual tuning refers to a set of HPO methods where practitioners choose hyperparameter configurations based on explicit and implicit knowledge and influences of contextual factors (e.g.,  findings presented in literature and the ML model architecture). Dependence of manual tuning on practitioners’ experiences and even unconscious rationales for decisions makes manual tuning very individual to practitioners, rendering the explication of manual tuning difficult. Thus, manual tuning is usually hardly replicable (Musgrave et al., 2020; Simon et al., 2023). Commonly, only intermediate information (e.g.,  used hyperparameter) is available to reproduce HPO methods in manual tuning, while reasons for selecting hyperparameter configurations often remain unclear. Because formalization of HPO methods in manual tuning is difficult, the number of strategies for manual tuning applied by practitioners is unknown. Common strategies include starting optimization from well-performing hyperparameter values (Xin et al., 2021), using a sound experiment design for incremental improvement (Godbole et al., 2023a), and eliminating irrelevant hyperparameters from the search space (Radosavovic et al., 2020). In addition, difficulty in explication make it hard to evaluate the sample efficiency of HPO methods in manual tuning. Although hard evidence that manual tuning generally outperforms advanced HPO methods (e.g., Bayesian optimization) is lacking, prior publications offer initial evidence that advanced HPO methods can outperform methods in manual tuning in several use cases (Feurer et al., 2016; Melis et al., 2018; Chen et al., 2018; Zhang et al., 2021; Kadra et al., 2021; Zöller et al., 2023).
@@
To tackle shortcomings of manual tuning, AutoML (Feurer and Hutter, 2019) is envisioned to automate all aspects related to development of ML models in a problem-agnostic manner, for example, by programmatic HPO methods. Research on AutoML has taken a foremost technical stance on HPO (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)). Typical works (e.g., (Bergstra et al., 2011; Jamieson and Talwalkar, 2016; Golovin et al., 2017; Falkner et al., 2018; Li et al., 2018a; He et al., 2018; Wang et al., 2021c)) focus on performance optimization of ML models in terms of smaller generalization errors, smaller ML model size, or lower latency. Research on AutoML investigates HPO from a mathematical perspective and often treats HPO as a black-box optimization problem: Given a problem instance in the form of a dataset and a loss function, a black-box optimizer (e.g., Bayesian optimization) searches for hyperparameter configurations in a predefined search space to enhance an ML model in terms of a given metric (e.g., accuracy on a validation set). Multi-objective optimization methods can be used to specify additional properties of the resulting ML model, such as low model complexity, fast inference, and algorithmic fairness (Gardner et al., 2019; Binder et al., 2020; Karl et al., 2023; Dooley et al., 2022).
@@
Although easy to implement, parallelize, and reproduce, grid search has become increasingly unsuited for modern HPO problems due to the curse of dimensionality (Bergstra and Bengio, 2012). In practice, not all hyperparameters have a similar impact on the final performance of ML models (Bergstra and Bengio, 2012; Rijn and Hutter, 2018). Due to the rigid search strategy, grid search tends to spend a large portion of the optimization budget on exploring irrelevant parts of the search space (Bergstra and Bengio, 2012). Put differently, sample efficiency of grid search tends to be lower compared to later HPO methods (e.g., random search, Bayesian optimization, and evolutionary algorithms (Snoek et al., 2012; Eggensperger et al., 2013; Turner et al., 2021)), in particular, because grid search cannot make use of the low effective dimensionality of HPO problems (Bergstra et al., 2011).
Random search refers to the process of sampling random hyperparameter configurations from a defined search space until a budget is exhausted (Bergstra and Bengio, 2012). Random search effectively deals with hyperparameters with different importance and reaches better sample efficiency in high-dimensional search spaces than grid search if some hyperparameters influence the performance of the ML model more than others (Bergstra and Bengio, 2012). Random search can be reproduced if the used search space, the randomness generator, and the corresponding seed are known.
To cope with large search spaces, the pure exploration by grid search and random search was complemented by the exploitation of knowledge of well-performing regions in sets of possible hyperparameter configurations (Snoek et al., 2012). An established approach for combining exploitation and exploration is evolutionary optimization. Inspired by biological evolution, evolutionary algorithms refer to processes of mutating a population of potential solutions iteratively to obtain solutions with better performance. Evolutionary algorithms often perform well in optimizing black-box functions (Olson and Moore, 2019) but are rather inefficient in terms of samples (Bischl et al., 2023). HPO based on evolutionary algorithms can be reproduced with fixed random seeds if the search space and randomness generator are known.
@@ 2.2. Research on Hyperparameter Optimization
Complementary to the technical focus on HPO, studies on human experiences of using programmatic HPO methods aim at better understanding practitioner perceptions of advantages and disadvantages of HPO methods (Gil et al., 2019; Wang et al., 2019b, 2021b). Practitioners acknowledge benefits of programmatic HPO methods, which are commonly related to faster turn-around time for building ML models and, thus, higher productivity in developing ML models (Wang et al., 2019b, 2021a). Automatically tuned ML models are often used by practitioners to create initial baseline ML models for subsequent manual tuning or to gain data insights (Wang et al., 2019b, 2021a; Crisan and Fiore-Gartland, 2021). However, AutoML practitioners often bemoan insufficient confidence in the results of programmatic HPO methods and, therefore, refuse to blindly use ML models optimized with programmatic HPO methods (Lee and Macke, 2020; Wang et al., 2019b; Drozdal et al., 2020; Khuat et al., 2023). Even though acknowledging that programmatic HPO is useful to develop well-performing ML models (Wang et al., 2021a), practitioners often refuse to use those HPO methods to not be accountable for ML models they do not understand (Drozdal et al., 2020). Missing confidence is mostly associated with the perceived black-box approach of programmatic HPO methods, such as Bayesian optimization and evolutionary algorithms, that inhibits sufficient transparency of the optimizer internals. Practitioners wish for augmentation of their daily data science work (e.g.,  through guidance) instead of fully automating it (Crisan and Fiore-Gartland, 2021).
Research on human-guided HPO focuses on involving humans in programmatic HPO methods to improve HPO with dormant domain expertise (Wang et al., 2019b; Higuchi et al., 2021; Wang et al., 2019a). This requires identifying how and when to involve humans in HPO to achieve the best combination of automation and human knowledge (Crisan and Fiore-Gartland, 2021). Especially involvement of practitioners in ML model development, including HPO, appears promising for a higher level of automation. For other tasks, including data acquisition and requirement analysis, practitioners prefer strong human involvement with a low level of automation (Wang et al., 2021b). Interactions of practitioners with software tools for programmatic HPO were structured into different modes of cooperation between practitioners and software tools, ranging from manual tuning to full automation, in the literature (Lee and Macke, 2020; Crisan and Fiore-Gartland, 2021; Wang et al., 2021b).
@@ 3. Methods
We applied a mixed-methods research approach incorporating two principal steps. First, we conducted semi-structured interviews with ML experts to develop a set of commonly used HPO methods, goals pursued by using HPO methods, and contextual factors that influence the choice for HPO methods. Second, we performed a survey using an online questionnaire to collect evidence for the external validity of the interviews. The following details the two steps.

### v4

@@ Abstract.
Programmatic hyperparameter optimization (HPO) methods, such as Bayesian optimization and evolutionary algorithms, are highly sample-efficient in identifying optimal hyperparameter configurations for machine learning (ML) models. However, practitioners frequently use less efficient methods, such as grid search, which can lead to under-optimized models. We suspect this behavior is driven by a range of practitioner-specific motives. Practitioner motives, however, still need to be clarified to enhance user-centered development of HPO tools. To uncover practitioner motives to use different HPO methods, we conducted 20 semi-structured interviews and an online survey with 49 ML experts. By presenting main goals (e.g.,  increase ML model understanding) and contextual factors affecting practitioners’ selection of HPO methods (e.g., available computer resources), this study offers a conceptual foundation to better understand why practitioners use different HPO methods, supporting development of more user-centered and context-adaptive HPO tools in automated ML.
@@ 1. Introduction
The performance of machine learning (ML)models is highly sensitive to their hyperparameter configurations (Bergstra and Bengio, 2012; Melis et al., 2018; Henderson et al., 2018; Chen et al., 2018; Zhang et al., 2021; Kadra et al., 2021). Identifying optimal hyperparameter configurations for training ML models is, however, a complex and often daunting task—even for seasoned ML experts—because of large search spaces of hyperparameter values and commonly unknown relationships between ML model performance, hyperparameter configuration, and dataset. Consequently, practitioners—individuals regularly involved in the development of viable ML models that are meant for productive use in research or industry—experiment with various hyperparameter configurations to identify optimal ones, often relying on trial and error. This iterative process of exploring, testing, and adjusting hyperparameter configurations related to ML models is known as hyperparameter optimization (HPO).
@@
HPO by manual tuning is often cumbersome, tedious, and error-prone. To help practitioners with HPO, research with a technical focus (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)) developed several programmatic HPO methods, including grid search, random search, Bayesian optimization, and evolutionary algorithms, and implemented such methods as software tools (e.g., Hyperopt and Hyperband; (Bergstra et al., 2015; Li et al., 2018a)). Existing HPO methods, programmatic and non-programmatic ones, differ considerably in the way they optimize hyperparameter configurations. Such differences lead some HPO methods to superiority over others. That superiority is usually demonstrated using conventional performance metrics from computer science, including minimization of generalization errors and increase of sample efficiency (Zöller and Huber, 2021; Gijsbers et al., 2022; Lindauer et al., 2022). However, practitioners often use HPO methods that are inferior according to conventional performance metrics (Bouthillier and Varoquaux, 2020). For example, practitioners often prefer to perform grid search over the more sample-efficient Bayesian optimization (Snoek et al., 2012).
@@
The dominant use of seemingly inferior HPO methods suggests that practitioners have motives beyond the fulfillment of conventional goals pursued in HPO, such as increasing ML model performance. However, the motives behind practitioners’ choices of HPO methods—shaped by contextual factors—remain unclear. This unclarity about why practitioners use which HPO methods inhibits more user-centered development of HPO methods and tools that support practitioners in attaining their goals, especially those beyond conventional performance metrics. To support development of more user-centered HPO methods and tools for automated ML (AutoML)(Vanschoren, 2019), practitioner motives for HPO need to be better understood. We approach the following research question: Why do practitioners choose different HPO methods?
@@
We applied a two-step research approach consisting of an interview study and a survey study based on an online questionnaire. First, we conducted semi-structured interviews with 20 ML experts to unveil HPO methods that practitioners commonly use, the goals pursued by practitioners when applying HPO methods, and the contextual factors that influence practitioners’ choices of HPO methods to attain goals. Second, we performed an online survey with 49 participants to collect evidence for the external validity of the relevance of the HPO methods, goals, and contextual factors identified in the interviews.
@@
Our main ambition is to support more user-centric development of automated ML (AutoML) methods and tools for HPO by bridging the gap between technical advancements and human factors. By integrating practitioner motives into HPO research, we aim to ensure that future tools are not only technically sound but also aligned with real-world needs and decision-making processes. In particular, this work has three main contributions. First, this work presents a conceptual foundation constituted of principal goals (e.g., improving ML model performance and target audience compliance) and contextual factors (e.g., compute resources and method traceability) influencing selections of HPO methods. This is useful to enhance human-in-the-loop ML by clarifying information needs and improving practitioner engagement. Moreover, the conceptual foundation supports goal-driven development of HPO methods and tools and new benchmarks with a focus on human factors. Second, we present a mapping of HPO methods, goals, and contextual factors to explain practitioners’ decision-making. This mapping informs user-centered HPO design and adaptive automation features. Moreover, it highlights key input parameters for developing context-sensitive HPO tools. Third, we analyzed practitioners’ perceived success with different HPO methods, revealing strengths and areas for improvement. These findings contribute to the development of more effective decision-support systems for HPO method selection. Moreover, the reported success rates help identify limitations of HPO methods, guiding development of better-tailored and more effective HPO methods and tools.
@@ 2. Background and Related Work
Driven largely by technical advancements, HPO is a key research area with significant potential to shape ML model development (Hutter et al., 2018). To support a better understanding of the results of this study, we briefly describe the technical foundations of principal HPO methods. Moreover, we elucidate related research on involving human beings in automated HPO, a prominent field within AutoML research.
@@ 2.1. Hyperparameter Optimization Methods
There are five principal, model-agnostic HPO methods: manual tuning, grid search, random search, evolutionary algorithms, and Bayesian optimization. [^1] There are also gradient-based methods, but because these methods are often perceived as being brittle and are often model-specific, they are out of the scope of this study. These HPO methods are briefly described below.
@@
Manual tuning refers to a set of HPO methods where practitioners choose hyperparameter configurations based on explicit and implicit knowledge and influences of contextual factors. The dependency of manual tuning on practitioners’ experiences and even unconscious heuristics in decisions (Gigerenzer and Brighton, 2009) makes manual tuning very individual to practitioners, rendering the explication of manual tuning difficult. Thus, manual tuning is usually hardly replicable (Musgrave et al., 2020; Simon et al., 2023). Commonly, only intermediate information (e.g.,  tuned hyperparameters) is available, while reasons for selecting hyperparameter configurations often remain unclear. This complicates the formalization of specific HPO methods in manual tuning, making it unclear how practitioners actually proceed. Common strategies include starting optimization from well-performing hyperparameter values (Xin et al., 2021), using a sound experiment design for incremental improvement (Godbole et al., 2023a), and removing irrelevant hyperparameters from the search space (Radosavovic et al., 2020). In addition, difficulty in explication make it hard to evaluate the sample efficiency of strategies in manual tuning. While evidence that manual tuning outperforms advanced HPO methods is lacking, prior publications offer initial evidence that advanced HPO methods can outperform manual tuning in several use cases (Feurer et al., 2016; Melis et al., 2018; Chen et al., 2018; Zhang et al., 2021; Kadra et al., 2021; Zöller et al., 2023).
@@
To tackle shortcomings of manual tuning, AutoML (Feurer and Hutter, 2019) is envisioned to automate all aspects related to development of ML models in a problem-agnostic manner, for example, by programmatic HPO methods. Research on AutoML has primarily approached HPO from a technical perspective (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)). Typical works (e.g., (Bergstra et al., 2011; Jamieson and Talwalkar, 2016; Golovin et al., 2017; Falkner et al., 2018; Li et al., 2018a; He et al., 2018; Wang et al., 2021c)) focus on performance optimization of ML models in terms of smaller generalization errors, smaller ML model size, or lower latency. Often, such works investigate HPO from a mathematical perspective and treat HPO as a black-box optimization problem: given a problem instance in the form of a dataset and a loss function, a black-box optimizer (e.g., Bayesian optimization) searches for hyperparameter configurations in a predefined search space to enhance an ML model in terms of a given metric (e.g., accuracy on a validation set). Multi-objective optimization methods can be used to specify additional properties of the resulting ML model, such as algorithmic fairness, fast inference, and low model complexity (Gardner et al., 2019; Binder et al., 2020; Karl et al., 2023; Dooley et al., 2022).
@@
Although easy to implement, parallelize, and reproduce, grid search has become increasingly unsuited for modern HPO problems due to the curse of dimensionality (Bergstra and Bengio, 2012). In practice, not all hyperparameters have a similar impact on the final performance of ML models (Bergstra and Bengio, 2012; Rijn and Hutter, 2018). Due to its rigid search strategy, grid search often allocates much of the optimization budget to less relevant regions of the search space (Bergstra and Bengio, 2012). Put differently, sample efficiency of grid search tends to be lower compared to later HPO methods (e.g., random search, Bayesian optimization, and evolutionary algorithms (Snoek et al., 2012; Eggensperger et al., 2013; Turner et al., 2021)), in particular, because grid search cannot make use of the low effective dimensionality of HPO problems (Bergstra et al., 2011).
Random search refers to the process of sampling random hyperparameter configurations from a defined search space until a budget is exhausted (Bergstra and Bengio, 2012). Random search handles hyperparameters of varying importance more effectively than grid search and achieves greater sample efficiency in high-dimensional search spaces, especially when hyperparameters have differing influence on ML model performance (Bergstra and Bengio, 2012). Random search can be reproduced if the used search space, the randomness generator, and the corresponding seed are known.
To address large search spaces, grid search and random search have been supplemented by methods that exploit knowledge of well-performing regions within the set of possible hyperparameter configurations (Snoek et al., 2012). A well-established approach for balancing exploration and exploitation is evolutionary optimization. Inspired by biological evolution, evolutionary algorithms iteratively mutate a population of candidate solutions to obtain solutions with better performance. Evolutionary algorithms often perform well in optimizing black-box functions (Olson and Moore, 2019) but are rather inefficient in terms of samples (Bischl et al., 2023). HPO based on evolutionary algorithms can be reproduced with fixed random seeds if the search space and randomness generator are known.
@@ 2.2. Research on Hyperparameter Optimization
Studies exploring practitioners’ experiences with programmatic HPO methods provide valuable insights into the perceived advantages and disadvantages of these methods (Gil et al., 2019; Wang et al., 2019b, 2021b). Practitioners acknowledge advantages of programmatic HPO methods, which are commonly related to faster turn-around time for building ML models and, thus, higher productivity in developing ML models (Wang et al., 2019b, 2021a). Automatically tuned ML models are often used by practitioners to create initial baseline ML models for subsequent manual tuning or to gain data insights (Wang et al., 2019b, 2021a; Crisan and Fiore-Gartland, 2021). However, AutoML practitioners often bemoan insufficient confidence in the results of programmatic HPO methods and, therefore, refuse to blindly use ML models optimized with programmatic HPO methods (Lee and Macke, 2020; Wang et al., 2019b; Drozdal et al., 2020; Khuat et al., 2023). Even though acknowledging that programmatic HPO is useful to develop well-performing ML models (Wang et al., 2021a), practitioners often refuse to use those HPO methods to not be accountable for ML models they do not understand (Drozdal et al., 2020). A lack of confidence is often linked to the perceived black-box nature of programmatic HPO methods, such as Bayesian optimization and evolutionary algorithms, limiting transparency of optimizer internals. Practitioners prefer support that augments their daily data science work (e.g., through guidance) rather than fully automating it (Crisan and Fiore-Gartland, 2021).
Research on human-guided HPO focuses on involving humans in programmatic HPO methods to improve HPO with dormant domain expertise (Wang et al., 2019b; Higuchi et al., 2021; Wang et al., 2019a). This requires identifying how and when to involve humans in HPO to achieve the best combination of automation and human knowledge (Crisan and Fiore-Gartland, 2021). Especially involvement of practitioners in ML model development, including HPO, seems promising for a higher level of automation. For other tasks, including data acquisition and requirement analysis, practitioners prefer strong human involvement with a low level of automation (Wang et al., 2021b). Interactions of practitioners with software tools for programmatic HPO were structured into different modes of cooperation between practitioners and software tools, ranging from manual tuning to full automation, in the literature (Lee and Macke, 2020; Crisan and Fiore-Gartland, 2021; Wang et al., 2021b).
@@ 3. Methods
We applied a mixed-methods research approach consisting of two main steps. First, we conducted semi-structured interviews with ML experts to develop a set of commonly used HPO methods, goals pursued by using HPO methods, and contextual factors that influence the choice for HPO methods. Second, we conducted a survey using an online questionnaire to collect evidence of the external validity of the interviews. The following details the two steps.